OpenStudy (anonymous):

The length of a rectangle is 4 inches less than twice the width of the rectangle. The perimeter of the rectangle is 58 inches. Write a system of equations to model this info.

OpenStudy (apoorvk):

Let the width of the triangle be 'x'. length = 4 less than twice the breadth = 2x - 4 perimeter of triangle = 2(length + breadth) = 58 .....{given} substitute length and breadth from above (in terms of 'x') in the equation for perimeter, and solve for 'x'.

OpenStudy (apoorvk):

You basically get: P = 2(2x-4 + x) = 58 can you solve this?